Differential Equations and Logarithmic-Exponential Series

David Marker

Van den Dries, Macintyre and I introduced the field of LE-series as a natural nonstandard model of the theory of the real field with exponentiation. The LE-series have a natural derivation and there seems to be a close relationship between solving differential equations in the LE-series and finding solutions in Hardy fields.

Address: Department of Mathematics, Statistics and Computer Science, Univesity of Illinois at Chicago, 851 S. Morgan St., Chicago IL 60607-7045, USA

E-mail: marker@math.uic.edu